# What Challenges Do Developers Face When Using Verification-Aware Programming Languages?

### v1

@@ V What improvements can be made to existing tools to overcome those challenges? (RQ3)
The user survey combined closed and open-ended questions. The closed questions aimed to validate the following specific proposals derived from our earlier analysis: 1) Create user-readable counter-examples to complement error messages. From our analysis, counter-examples not always seem helpful and their use was even discouraged in some cases; 2) Develop official tutorials on common verification exercises (e.g., Fibonacci sequence, greatest common divisor). Neither Dafny nor Frama-C provides guided tutorials on their official pages; 3) Implement a search feature in the documentation. Neither Dafny nor Frama-C include a search feature, and Frama-C’s documentation is only available in PDF format. Dafny’s online documentation lacks a persistent table of contents, unlike popular tools like Jest [^9] https://jestjs.io/docs/getting-started [last access 16th July 2024], and Pytest [^10] https://docs.pytest.org/en/8.2.x/contents.html [last access 16th July 2024]; 4) Expand the themes covered in the documentation based on the gaps identified in our analysis; and 5) Promote VA languages in mandatory courses, as these languages often appear only in optional courses, unlike testing, which is more frequently taught.
@@
Open-ended Questions. We identified 13 categories of improvement for VA languages (see Table V). The most frequently mentioned area was improvements in documentation and learning materials, suggested by five participants. These requests included better tutorials, guides, and documentation to assist both newcomers and experienced developers in using VA languages effectively. For example, one beginner developer, P28, noted the need for “more documentation on coming up with invariants.” Four participants mentioned improved usability, and three better education. When facing errors, two beginners and one experienced developer mentioned that better debugging would help them understand why the verification fails. P23 mentioned directly that “To be really useful in practice, a verification-aware language needs to be familiar, easy to write proofs in, predictable in terms of what can be verified, well-integrated with common tools, interoperable with other languages, and able to generate efficient code.” (P23). This aligns with the broader theme of integration with mainstream languages, mentioned by four participants.
Another suggestion by two participants is related to reusable blocks, P19 proposed that tools should “provide more automation to help ‘easy tasks’, so the programmer/prover can focus on complex tasks”, and P26, mentioned directly they would like to see “reusable bricks for common verification patterns.“ Three participants advocated for increased investment in verification tools, suggesting that funding and institutional support are necessary for widespread adoption. While the majority of participants maintained a positive outlook, skepticism regarding practical adoption was expressed by two participants. For instance, P13 remarked, “verification is expensive, and tends not to produce immediate value - something which corporate doesn’t like very much.”.
@@ VI Related Work
To the best of our knowledge, this is the first work to address challenges faced by developers using VA languages. In this section, we focus on related work on analysis and topic modelling of developer discussions, and on studies on challenges faced when using different programming languages.
@@
Developer Discussion Analysis and Topic Modelling. Prior studies have applied topic modelling methodologies to analyse developer discussions on platforms like Stack Overflow and GitHub, focusing on areas different from VA languages [9]. Barua et al. [7] analysed posts from a Stack Overflow data dump (available since 2014 on the Internet Archive) to understand the main general discussion topics on the platform. Bagherzadeh and Khatchadourian [3] used the official Stack Overflow data dump to study the topics of questions specifically about big data. Han et al. [25] studied what issues deep learning framework practitioners discuss, gathering data from Stack Overflow (official data dumps) and GitHub (official API). Abdellatif et al. [1] used the official Stack Overflow data dump as the dataset to identify the topics chatbot developers were asking about. They analysed heuristics to discover relevant tags, notably using only post titles for training data, claiming it was representative and reduced noise. Openja et al. [38] analysed Stack Overflow questions to understand topics discussed around Release Engineering, using the SOTorrent dataset created by Baltes et al. [4]. They performed topic extraction using significance and relevance heuristics for tags. El aoun et al. [17] investigated challenges in quantum software engineering using data gathered from GitHub and the Stack Exchange Data Explorer (for Stack Overflow). Al Alamin et al. [2] conducted an empirical study of developer discussions on low-code software development challenges. Peruma et al. [41] used the SOTorrent dataset to study the topics around Software Refactoring being asked by developers. They filtered posts to ensure relevance. These studies collectively demonstrate a well-established methodology for using topic modelling, primarily on online forum data, to explore and understand developer discussions across various software engineering areas. Our work builds on these previous studies used to analyse the challenges of different areas and applies them to the specific context of VA languages.
@@
Programming Language Challenges. Other studies have examined the challenges developers face when using programming languages, suggesting that even experienced developers struggle to learn new programming languages [46]. Ko et al. [31] identified six learning barriers in programming systems, which include understanding, design, use, and information barriers. These barriers align closely with the challenges reported by VA language users in our study, such as difficulty interpreting tool feedback and expressing proof strategies. Other studies focused on particular languages. For instance, Zhu et al. [53] focused on Rust, which uses compile-time checks to ensure memory safety through mechanisms like ownership and lifetimes. They analyzed 15,000 Stack Overflow posts and surveyed 101 Rust developers to find that Rust safety features impose a steep learning curve and that compiler messages often lack actionable information. This is in line with other work that identified onboarding and usability issues as barriers to the adoption of languages [22, 29]. Chakraborty et al. [10] examined the support ecosystem for newer languages such as Go, Swift, and Rust by analyzing both Stack Overflow and GitHub data. They identified gaps in documentation, which often drove developers to rely on community Q&A sites. Finally, Christakis and Bird [11] studied developer expectations regarding program analysis tools by surveying engineers at Microsoft. They found that high false-positive rates, poor tool integration, and unclear feedback were the main reasons developers rejected static analysis tools in practice. These findings echo the feedback-related problems reported by VA language users in our study and suggest that if tools fail to align with developers’ workflows or generate opaque error messages, they are unlikely to be adopted.
@@ VII Conclusion
This work analyzes the challenges developers face with VA languages, incorporating feedback from both beginners and experienced users. We trained an ML model on 1,420 questions, with 145 analyzed post-clustering. From this, 24 distinct challenges were identified and rated by 31 study participants Additionally, five improvement suggestions were evaluated, and 12 more were uncovered via open-ended responses.
This work can pave the way for future investigations of challenges and improvements, while also offering valuable insights to guide the design of new VA languages or the enhancement of existing ones to support broader adoption.
@@
Future Work. While this work answers the main research questions, several areas can be expanded. Future work includes expanding the dataset with further sources and carry out a large-size user study. In addition, newly identified challenges should be ranked based on relevance, in order to direct development towards support that can have a greater impact.

### v2

@@ IV What is the context of the first contact with Verification-Aware languages? (RQ2)
These results suggest that beginners usually first encounter VA languages at university, while experienced developers do so in professional settings and increasingly through self-teaching. These insights highlight the need for differentiated learning resources: universities should integrate VA languages into curricula for early exposure, while practitioners could benefit from self-guided resources and community support.
@@ V What improvements can be made to existing tools to overcome those challenges? (RQ3)
The user survey combined closed and open-ended questions. The closed questions aimed to validate the following specific proposals derived from our earlier analysis: 1) Create user-readable counter-examples to complement error messages. From our analysis, counter-examples do not always seem helpful and their use was even discouraged in some cases; 2) Develop official tutorials on common verification exercises (e.g., Fibonacci sequence, greatest common divisor). Neither Dafny nor Frama-C provides guided tutorials on their official pages; 3) Implement a search feature in the documentation. Neither Dafny nor Frama-C include a search feature, and Frama-C’s documentation is only available in PDF format. Dafny’s online documentation lacks a persistent table of contents, unlike popular tools like Jest, [^10] https://jestjs.io/docs/getting-started [last access 16th July 2024] and Pytest; [^11] https://docs.pytest.org/en/8.2.x/contents.html [last access 16th July 2024] 4) Expand the themes covered in the documentation based on the gaps identified in our analysis; and 5) Promote VA languages in mandatory courses, as these languages often appear only in optional courses, unlike testing, which is frequently taught.
@@
Open-ended Questions. We identified 13 categories of improvement for VA languages (see Table V). The most frequently mentioned area was improvements in documentation and learning materials, suggested by five participants. These requests included better tutorials, guides, and documentation to assist both newcomers and experienced developers in using VA languages effectively. For example, one beginner developer, P28, noted the need for “more documentation on coming up with invariants”. Four participants mentioned improved usability, and three better education. When facing errors, two beginners and one experienced developer mentioned that better debugging would help them understand why the verification fails. P23 mentioned directly that “To be really useful in practice, a verification-aware language needs to be familiar, easy to write proofs in, predictable in terms of what can be verified, well-integrated with common tools, interoperable with other languages, and able to generate efficient code.” (P23). This aligns with the broader theme of integration with mainstream languages, mentioned by four participants.
Another suggestion by two participants is related to reusable blocks, P19 proposed that tools should “provide more automation to help ‘easy tasks’, so the programmer/prover can focus on complex tasks”, and P26, mentioned directly they would like to see “reusable bricks for common verification patterns“. Three participants advocated for increased investment in verification tools, suggesting that funding and institutional support are necessary for widespread adoption. While the majority of participants maintained a positive outlook, skepticism regarding practical adoption was expressed by two participants. For instance, P13 remarked, “verification is expensive, and tends not to produce immediate value, something which corporate doesn’t like very much.”
@@ VI Related Work
To the best of our knowledge, this is the first work to address challenges faced by developers using VA languages. In this section, we focus on related work on analysis and topic modeling of developer discussions, and on studies on challenges faced when using different programming languages.
@@
Developer Discussion Analysis and Topic Modeling. Prior studies have applied topic modeling to analyse developer discussions on platforms like Stack Overflow and GitHub, focusing on areas different from VA languages [10]. Barua et al. [8] analysed posts from a Stack Overflow data dump to understand the main general discussion topics on the platform. Bagherzadeh and Khatchadourian [4] used the official Stack Overflow data dump to study the topics of questions about big data. Han et al. [25] studied what issues deep learning framework practitioners discuss, gathering data from Stack Overflow and GitHub. Abdellatif et al. [1] used the official Stack Overflow data dump to identify the topics chatbot developers were asking about. They analysed heuristics to discover relevant tags, using only post titles for training data, claiming it was representative and reduced noise. El aoun et al. [18] investigated challenges in quantum software engineering using data from GitHub and the Stack Exchange Data Explorer (for Stack Overflow). Al Alamin et al. [3] conducted an empirical study of developer discussions on low-code software development challenges. Peruma et al. [40] used the SOTorrent dataset to study the topics around Software Refactoring being asked by developers. These studies demonstrate a well-established methodology for using topic modeling, primarily on online forum data, to study developer discussions across various software engineering areas. Our work builds on these studies and applies them to the specific context of VA languages.
@@
Programming Language Challenges. Other studies have examined the challenges developers face when using programming languages, suggesting that even experienced developers struggle to learn new programming languages [44]. Ko et al. [31] identified six learning barriers in programming systems, which include understanding, design, use, and information barriers. These barriers align closely with the challenges reported by VA language users in our study, such as difficulty interpreting tool feedback and expressing proof strategies. Other studies focused on particular languages. For instance, Zhu et al. [51] focused on Rust. They analyzed 15,000 Stack Overflow posts and surveyed 101 Rust developers to find that Rust safety features impose a steep learning curve and that compiler messages often lack actionable information. This is in line with other work that identified onboarding and usability issues as barriers to the adoption of languages [22, 29]. Chakraborty et al. [11] examined the support ecosystem for newer languages such as Go, Swift, and Rust by analyzing both Stack Overflow and GitHub data. They identified gaps in documentation, which often drove developers to rely on community Q&A sites. Finally, Christakis and Bird [12] studied developer expectations regarding program analysis tools by surveying engineers at Microsoft. They found that high false-positive rates, poor tool integration, and unclear feedback were the main reasons developers rejected static analysis tools in practice. These findings echo the feedback-related problems reported by VA language users in our study and suggest that if tools fail to align with developers’ workflows or generate opaque error messages, they are unlikely to be adopted.
@@ VII Conclusion
This work analyzes the challenges developers face with VA languages, incorporating feedback from both beginners and experienced users. We trained an ML model on 1,420 questions, with 145 analyzed post-clustering. From this, 24 distinct challenges were identified and rated by 31 study participants. Additionally, five improvement suggestions were evaluated, and 12 more were uncovered via open-ended responses.
Our findings suggest that VA language adoption is primarily blocked by lack of documentation, and usability and ecosystem limitations, highlighting key areas for improvement. For example, a challenge across all experience levels lies in designing and completing proof strategies, underscoring the need for better tool support and guidance. Additionally, the mismatch found between online discussions and reported challenges (e.g., frame specifications and proof termination clauses, were identified in Stack Overflow, but not confirmed as challenges), points to the importance of empirical user studies, as research based solely on online discussions may not fully capture the real experience of practitioners (or vice-versa).
This work can pave the way for future research on challenges and improvements, while also offering valuable insights to guide the design and/or enhancement of VA languages to support broader adoption. Future work includes expanding the dataset with further sources and carry out a large-size user study. In addition, newly identified challenges should be ranked based on relevance, in order to direct development towards support that can have a greater impact.
@@
Acknowledgments
Alexandra Mendes was supported by an Amazon Research Award, Fall 2024. Carolina Carreira was financed by National Funds through the FCT - Fundação para a Ciência e a Tecnologia, I.P. (Portuguese Foundation for Science and Technology) within the project VeriFixer, with reference 2023.15557.PEX (DOI: 10.54499/2023.15557.PEX), and by funding from FCT under grant PRT/BD/153739/2021. We thank João F. Ferreira for his valuable comments on this work.